The time spent waiting in the line is approximately normally distributed. The mean waiting time is 5 minutes and the variance of the waiting time is 4. Find the probability that a person will wait for between 6 and 9 minutes. Round your answer to four decimal places
Answer #1

Solution :

Given that ,

mean = = 5

2 =4

standard deviation = = 2  

P(6< x < 9) = P[(6-5) /2 < (x - ) / < (9-5) / 2)]

= P( 0.5< Z <2 )

= P(Z < 2) - P(Z < 0.5)

Using z table   

= 0.9772-0.6915

probability= 0.2857